OpenStudy (lastdaywork):

Resolution of vectors: Let A, B, C be non-coplanar 'unit' vectors and a, b, c be scalars such that - rR = aA + bB + cC where r is a scalar and R is a 'unit' vector Given that R makes an angle of α, β, γ with A, B, C respectively ; find the value of a, b, c in terms of r.

OpenStudy (ikram002p):

(cos α )^2+(cos β )^2+(cos γ )^2=1 cos α=rR.A/a somthing like this

OpenStudy (lastdaywork):

(cos α )^2+(cos β )^2+(cos γ )^2=1 I think this equation is valid only for rectangular co-ordinates (and I am working with skew)

OpenStudy (ikram002p):

forget the equation use angle of direction law

OpenStudy (lastdaywork):

I have solved the same for 2D system  r sin(α) = b sin(α+β) Similarly, r sin(β) = a sin(α+β) But I can extend the same to 3D system :(

OpenStudy (lastdaywork):

*can't

OpenStudy (lastdaywork):

BTW "what's angle of direction law" ??

OpenStudy (ikram002p):

u knw angle btw i ,j,k r called angle of direction

OpenStudy (ikram002p):

btw any vector+i,j,k

OpenStudy (ikram002p):

for a vector v=<v1,v2,v3> direction angels α, β, γ cos α=v./||v|| ||i|| and so on

OpenStudy (ikram002p):

v.i sry not v alone

OpenStudy (ikram002p):

cos α=v.i/||v|| ||i|| cos β =v.j/||v|| ||j|| cos γ=v.k/||v|| ||k||

OpenStudy (lastdaywork):

But I don't have the value of v.i ; v.j ; v.k

OpenStudy (ikram002p):

:|

OpenStudy (ikram002p):

lol its only a hint

OpenStudy (ikram002p):

A,B,C unit vector use them insted of i,j,k rR is ur v

OpenStudy (lastdaywork):

There's a small problem; for rectangular co-ordinate system ; we can easily say v.i = v1 But in this case; v.i = r a cos(α) v.i = a + b(B.A) + c(C.A) Now I don't know the angle b/w B and A so I cant solve (B.A) ; same goes for (C.A)

OpenStudy (ikram002p):

no it not a problem cuz in ur qs the definition of α, β, γ is for ABC not i j k

OpenStudy (ikram002p):

i guess it will be like this cos α=a/r cos β =b/r cos γ=c/r

OpenStudy (lastdaywork):

Oh..I should have written it like this rR.A = r a cos(α) rR.A = a + b(B.A) + c(C.A) Now can you see where am I stuck ?

OpenStudy (lastdaywork):

Sorry, rR.A = r cos(α) rR.A = a + b(B.A) + c(C.A) So cos(α) = ( a + b(B.A) + c(C.A) ) / r

OpenStudy (ikram002p):

:| idk im confused with this too ill try it ltr

OpenStudy (lastdaywork):

Thanks for your time :)

OpenStudy (ikram002p):

np lol , its v intresting qs :o

OpenStudy (ikram002p):

try this book for vecor calcules Calculus.Early.Transcendentals.10th.edition.by.Stephen.Davis.Howard.Anton.and.Irl.Bivens very helpfull

OpenStudy (lastdaywork):

Thanxx ^_^

OpenStudy (lastdaywork):

Just to clarify...the problem is not yet solved XD

OpenStudy (lastdaywork):

The formula will finally lead to cos(α) = ( a + b(B.A) + c(C.A) ) / r in which IDK the value of (B.A) and (C.A)

OpenStudy (lastdaywork):

Projection of R on B may not be equal to b We can resolve aA into components along Cartesian co-ordinates but (I think) the final answer should be independent of the choice of coordinate system.

OpenStudy (lastdaywork):

It's not a textbook Q..I just wanted to find the general formula for resolving a 3D vector along any three non-coplanar vectors.

OpenStudy (lastdaywork):

Notation Small letters denote magnitude Capital letters denote Unit Vectors We have r R = a A + b B + c C We are given the value of r and the angles R makes with A, B, C. I need to find the value of a, b, c I have already solved this for 2-dimension (see 4th post from top) ; but can't extend the solution to 3-dimension.

OpenStudy (lastdaywork):

In your notations ; I need to find the magnitude of vector a in terms of magnitude of resultant vector.